# GroundingSuite: Measuring Complex Multi-Granular Pixel Grounding

### v2

@@ 1 Introduction
Table 1: Comparisons with previous Referring Expression Segmentation benchmark. ‘Cat.’ and ‘Len.’ denote the number of categories and average text length; Stuff: includes stuff classes; Part: includes part-level annotations; Multi: supports multi-object references; Single: supports single-object references;
@@ 4.2 Human Data Curation
To ensure the quality of  GSEval, we involve human reviewers for manual checks. Specifically, human reviewers carefully select and verify images across four distinct categories:
  • stuff segmentation: 1000 images featuring background elements requiring context-aware localization, such as ‘sky’ or ‘sea’;
  • part-level segmentation: 500 images focusing on object components demanding fine-grained understanding, such as ‘camera on the phone’ or ‘man’s beard’;
  • multi-object segmentation: 800 images containing complex referential relationships between multiple entities, such as ‘flock of sheep’ or ‘two dogs’;
  • single-object segmentation: 1500 images showcasing diverse appearance variations, such as ‘a brown cat’ or ‘a colorful parrot’.
@@ 5.3 Benchmark Results on GSEval-BBox
Gemini-1.5-Pro [26] Proprietary Model - 86.7 58.5 61.7 77.3 74.3
Doubao-1.5-vision-pro [25] - 80.0 28.8 85.2 53.3 64.2
@@
InternVL2.5-78B [5] Open-source Model 92.3 85.3 16.8 63.2 55.7 62.2
@@ 6.2 Comparisons with Other Automated Datasets
For fair comparisons, we randomly selected 100k samples from both GranD and the proposed GSTrain-10M to train models under the EVF-SAM and LISA framework.
@@ Results.
The result in Tab. 5 demonstrates that our GSTrain dataset consistently outperforms GranD across all benchmarks. With EVF-SAM, GSTrain improves performance by 2.8, 0.9, and 1.2 percentage points on RefCOCO, gRefCOCO, and RefCOCOm respectively. The gains are even more significant with LISA-7B, showing a 2.5 point improvement on RefCOCO and a remarkable 8.0 point gain on RefCOCOm (37.6 vs. 29.6), indicating GSTrain’s particular effectiveness with complex linguistic expressions.
@@ 7 Conclusion
Appendix A GSEval Details
@@
Figure 5: The word cloud of GSEval
Figure 6: More selected samples from our GSEval. Stuff class and part level
Figure 7: More selected samples from our GSEval. Multi object and single object

### v3

@@ 1 Introduction
Table 1: Comparisons with previous Referring Expression Segmentation benchmark. ‘Cat.’ and ‘Len.’ denote the number of categories and average text length; Stuff: includes stuff classes; Part: includes part-level annotations; Multi: supports multi-object references; Single: supports single-object references.
@@ 4.2 Human Data Curation
To ensure the quality of  GSEval, every image was manually selected and verified by human reviewers. This process resulted in a dataset organized into four distinct categories:
  • Stuff Segmentation: Comprises 1,000 images of background elements that require context-aware localization, such as sky and sea.
  • Part-level Segmentation: Consists of 500 images that demand fine-grained understanding of object components, such as the camera on a phone or a man’s beard.
  • Multi-object Segmentation: Includes 800 images with complex referential relationships between entities, like a flock of sheep or two dogs.
  • Single-object Segmentation: Contains 1,500 images that showcase diverse object appearances, such as a brown cat and a colorful parrot.
@@ 5.1 Benchmark Details
Figure 3: Visualizations of samples from GSEval (Stuff & Part). We showcase some examples with stuff and part annotations from GSEval.
Figure 4: Visualizations of samples from GSEval (Single & Multi). We showcase some segmentation annotations of single and multiple objects from GSEval.
@@ 5.3 Benchmark Results on GSEval-BBox
Gemini-1.5-Pro [27] Proprietary
Model
- 86.7 58.5 61.7 77.3 74.3
Doubao-1.5-thinking-vision-pro[25] 91.3 86.8 69.7 83.1 74.3 79.0
Doubao-1.5-vision-pro [26] 91.6 80.0 28.8 85.2 53.3 64.2
@@
InternVL3-78B[41] Open-source
Model
91.4 69.3 13.8 61.1 49.5 52.9
InternVL3-8B[41] 89.6 77.3 8.4 56.6 46.5 52.3
InternVL2.5-78B [5] 92.3 85.3 16.8 63.2 55.7 62.2
@@ 6.2 Comparisons with Other Automated Datasets
For fair comparisons, we randomly selected 100k samples from both GranD and the proposed GSTrain-10M to train models under the LAVT, EVF-SAM and LISA framework.
@@ Results.
As evidenced in Tab. 5, our proposed GSTrain dataset consistently outperforms the GranD dataset across all evaluated models and benchmarks. This benefit is clear for the specialist model, LAVT, which sees performance boosts of 2.2, 1.1, and 2.0 percentage points on RefCOCO, gRefCOCO, and RefCOCOm, respectively. The advantages of GSTrain are even more pronounced for the MLLM-based models. EVF-SAM, for instance, achieves gains of 2.8, 0.9, and 1.2 points across the same benchmarks. The most significant improvement is observed with LISA-7B, which not only improves by 2.5 points on RefCOCO but also exhibits a substantial 8.0-point increase on RefCOCOm (37.6 vs. 29.6). This considerable performance gain on a benchmark known for its linguistic complexity underscores that GSTrain provides the rich, high-quality data crucial for enabling advanced models to better interpret fine-grained textual descriptions.
@@
Expression Length $\leq$10 11 - 15 $\geq$16
w/o GSTrain-10M 79.2 77.4 81.7
w/   GSTrain-10M 79.7
+0.5
78.0
+0.6
82.8
+1.1
Table 6: Performance comparison on RefCOCOg across different referring expression lengths.
Filter Type RefCOCO gRefCOCO RefCOCOm
No Filter 36.9 26.2 25.9
CLIP-based Filter 60.7 28.0 31.5
IoU-based Filter (Ours.) 63.5
+2.8
28.9
+0.9
32.7
+1.2
Table 7: Ablation study on the effectiveness of different filter types.
6.3 Analysis of long expressions
Our quantitative results (Tab. 6) demonstrate that training the EVF-SAM model with our GSTrain-10M dataset yields a significant 1.1 cIoU point improvement on RefCOCOg for expressions of 16 words or more. We attribute this to the textual diversity and complexity of GSTrain-10M. More samples are visualized in revision.
6.4 Analysis of different filter types
We compared different filtering methods: no filter (baseline for unfiltered data), a CLIP-based filter (GranD), and our proposed IoU-based filter. The results in Tab. 7 indicate that filtering is essential for generating high-quality data. Significantly, our IoU-based filter demonstrates superior performance over the CLIP-based approach, achieving a more substantial reduction in labeling noise and enhancing data precision.
@@ 7 Conclusion
Appendix A Dataset Details
@@
Figure 7: The word cloud of GSEval.
Figure 8: Distribution of mask sizes of GroundingSuite (zoom in for more detailed information).
Figure 9: Distribution of mask positions of GroundingSuite (zoom in for more detailed information).
Figure 10: Partial statistics of fine-grained class IoU (zoom in for more detailed information).
In addition, we analyze mask size distribution (Fig. 8) and position distribution (Fig. 9) towards understanding mask size and position biases, which demonstrate GSTrain (Sampled from SA-1B) and GSEval (Sampled from unlabeled COCO) have significantly different distributions. We add a detailed breakdown of performance for different object classes. Partial statistics are presented in Fig. 10.